OpenStudy (anonymous):

Two AP contains the same no. of terms. The ratio of the last term of the first AP to the first term of second AP is equal to the ratio of the last term of second AP to the first term of the first AP and is equal to 4. The ratio of the sum of the first AP to that of the second AP is 2. Find the ratio of the differences of the progressions.

OpenStudy (anonymous):

Let the terms be n: So they contain the same terms: n Last term of first AP : \(a_n = a_x + (n-1)d_x\) So: \[\frac{a_x+ (n-1)d_x}{a_y} = \frac{a_y + (n-1)d_y}{a_x} = 4\] \[\large \frac{\frac{n}{2}(2 a_x + (n-1)d_x)}{\frac{n}{2}(2 a_y + (n-1)d_y)} = 2\]

OpenStudy (anonymous):

\(a_x\) : first term of first AP \(a_y\) : first term of second AP.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Wait let me find easier method to solve this..

OpenStudy (anonymous):

This has lengthy calculations.. But still I am close to it..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think today is not my day..

OpenStudy (anonymous):

@mukushla can help.

OpenStudy (anonymous):

\[a_x + (n-1)d_x = 4a_y\] \[a_y + (n-1)d_y = 4a_x\]

OpenStudy (anonymous):

\[\frac{a_x + 4a_y}{a_y + 4a_x} = 2 \implies \frac{a_x}{a_y} = \frac{2}{7}\]

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

Got it I think..

OpenStudy (anonymous):

Find \((n-1)\) from the two equations...

OpenStudy (anonymous):

\[n-1 = \frac{4a_y - a_x}{d_x}\] \[n - 1 = \frac{4a_x - a_y}{d_y}\]

OpenStudy (anonymous):

Equate them..

OpenStudy (anonymous):

\[\frac{d_x}{d_y} = \frac{4a_y - a_x}{4a_x - a_y}\] And you know : \(a_x = \frac{2}{7}a_y\) Try to go further..